OpenStudy (anonymous):

Determine the coordinates of all the points of intersection of: 1) y = x^2 +3x - 4 and y = 5x +11 2) y = cos x and y = sin x in quadrant 1

OpenStudy (anonymous):

put 5x+11 in 1st eqn and get 5 and -3 as ur x's points then solve for y for both values and get 36 and -4 as ur values at x =45 degrees its a std alue

OpenStudy (chaise):

y = 5x+11 gradient = 5 y intercept is = 11. x intersept is -11/5 I can help you with your parabola. Give me a second.

OpenStudy (chaise):

Do you want the solutions or do you want help achieving the solutions?

OpenStudy (anonymous):

y = x^2 +3x - 4 and y = 5x +11 Equate and solve for x 5x +11 = x^2 +3x -4 0=x^2 -2x -15 0 =(x^2 -5x +3x -15) 0=(x-5) (x +3) Solving we get x =5 and x = -3 Plug in to find y coordinate of the points y = 5x +11 x =5, y = 5*5 +11 = 25 +11 = 36 x =-3, y = 5*(-3) +11 = -15 +11 = -4 Points of intersection are (5,36) and (-3,-4)